What is the linear equation for 4x-2y=-1
Answer by fractalier(6550)

What you show, 4x-2y=-1, is a linear equation in standard form. If you need it in slope-intercept form, the most useful form for graphing, we solve for y...
4x - 2y = -1
-2y = -4x - 1
2y = 4x + 1
y = 2x + 1/2